OpenStudy (anonymous):

Which of the following is the simplified version of: I'm gonna draw it hold on

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } \log_{3} x+ 2 \log_{3} y +4 \log_{3} z\]

OpenStudy (anonymous):

Help? :(

OpenStudy (anonymous):

I can give you the options too?

OpenStudy (anonymous):

they have base 3? everyone?

OpenStudy (anonymous):

so u just need to add 4+2 + (1/2)

OpenStudy (anonymous):

yeah base 3

OpenStudy (anonymous):

add

OpenStudy (anonymous):

and u will get 6+1/2

OpenStudy (anonymous):

= 13/2

OpenStudy (anonymous):

ohh sorry i m doing it wrong

OpenStudy (anonymous):

hoa i hope u have a solution because i dont know what to do there

OpenStudy (anonymous):

ok, let me try

OpenStudy (anonymous):

= \[\log_{3}x^1/2 + \log_{3} y^2 +\log_{3}z^4 \]

OpenStudy (anonymous):

=\[\log_{3}\sqrt{x}y^2z^4 \]

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

to not be confused, re write as \[\log_{3} y^2 z^4 \sqrt{x}\]

OpenStudy (anonymous):

yes u r right